# TU Eindhoven, Eindhoven, the Netherlands

### v1

@@ 1 Introduction
Active inference agents are based on free energy functionals that rank policies on explorative and goal-directed behaviour [5, 7, 6, 16]. The expected free energy functional can be understood through its decomposition into a cross-entropy term between states and observations given action (“ambiguity”), and a Kullback-Leibler divergence between the posterior predictive and a goal prior distribution (“risk”) [7, 18, 4]. We show that Gaussian approximations of a non-linear observation function that are itself linear in the covariance matrix, e.g., first-order Taylor and the unscented transform [9], lead to ambiguity terms that are constant over states. This echoes an earlier finding that agents with a linear Gaussian state-space model exhibit a constant ambiguity term [10]. However, utilizing a second-order Taylor approximation induces a non-constant ambiguity term. Under this model, the agent will avoid states where the non-linear measurement function curves strongly. Our contributions are:
@@ 2 Problem statement
We want to plan a trajectory for a robot across a plane. The robot’s state at time $k$ is its planar position and time derivatives, $x_{k}\in\mathbb{R}^{D_{x}}$. The robot does not sense position directly, but has to infer it from noisy measurements $y_{k}\in\mathbb{R}^{D_{y}}$, produced by a sensor through a non-linear mapping $g:\mathbb{R}^{D_{x}}\rightarrow\mathbb{R}^{D_{y}}$ and measurement noise $v_{k}\in\mathbb{R}^{D_{y}}$. It accepts control inputs $u_{k}\in\mathbb{R}^{D_{u}}$ and moves according to linear dynamics with a transition matrix $A\in\mathbb{R}^{D_{x}\times D_{x}}$, control matrix $B\in\mathbb{R}^{D_{x}\times D_{u}}$ and process noise $e_{k}\in\mathbb{R}^{D_{x}}$. Overall, we consider robot systems described with discrete-time state-space models of the form:
@@
where $Q\in\mathbb{R}^{D_{x}\times D_{x}}_{+}$, $R\in\mathbb{R}^{D_{y}\times D_{y}}_{+}$ are noise covariance matrices.
@@ 2.0.1 Example
$g(x_{k})=\begin{bmatrix}\phi_{k}\\ d_{k}\end{bmatrix}=\begin{bmatrix}\sqrt{x_{1k}^{2}+x_{2k}^{2}}\\ \arctan(x_{1k},x_{2k})\end{bmatrix}\,,\qquad R=\begin{bmatrix}\rho^{2}_{1}&0\\ 0&\rho^{2}_{2}\end{bmatrix}\,,$ (4)
@@ 3.1 Probabilistic Model
This distribution is linear in $x_{k}$, and will allow for exact Bayesian filtering. But note that the parameters $\mu_{k}$, $\Gamma_{k}$ and $\Sigma_{k}$ may be nonlinear functions of $x_{k}$, depending on the type of Gaussian approximation (specifics treated in Section 4), and may be a richer representation of the effect of $g(x_{k})$.
@@ 3.3 Inferring controls
$q(y_{t}\>|\>x_{t})\triangleq p(y_{t}\>|\>x_{t})\,,\qquad q(x_{t},u_{t})\triangleq p(x_{t}\>|\>u_{t};\mathcal{D}_{k})q(u_{t})\,.$ (14)
@@
“Risk” refers to the Kullback-Leibler (KL) divergence between predicted and desired future observations. The inner integral in the risk term leads to a Gaussian distribution (Eq. 24) and the KL divergence between Gaussians is [3]:
@@
where $\Psi_{*}\triangleq(\mu_{*}-\mu_{t}\big{)}\big{(}\mu_{*}-\mu_{t})^{\intercal}$. “Ambiguity” refers to the conditional entropy of the future observations given the future states.
@@
$\mathcal{J}_{k}(u_{t})=\text{constants}+\frac{1}{2}\text{tr}\big{(}\Sigma_{*}^{-1}(\Sigma_{t}+\Psi_{*})\big{)}+\frac{1}{2}\ln\frac{\big{|}\Sigma_{t}-\Gamma^{\intercal}_{t}\bar{S}_{t}^{\text{-}1}\Gamma_{t}\big{|}}{\big{|}\Sigma_{t}\big{|}}\,.$ (29)
Note that $\Gamma_{t}$, $\Sigma_{t}$ and $\Psi_{*}$ depend on $u_{t}$. The above steps can be generalized to a longer time horizon $t=k+1,\dots k+T$. The prior is independent over time, so the joint factorizes as: $p(\bar{u})=\prod_{t=1}^{T}p(u_{t})$ (see Eq. 6). This means that the expected free energy function over $\bar{u}_{k}$ also factorizes to a sum of recursive expected free energy functions $\mathcal{J}_{k}(\bar{u}_{k})=\sum_{t=1}^{T}\mathcal{J}_{k}(u_{t})$, where the predicted state distribution parameters $\bar{m}_{t},\bar{S}_{t}$ are updated through the state transition (Eq. 10).
@@ 4 Gaussian approximations
$\mathbb{E}_{p(y_{t},x_{t}\>|\>u_{t};\mathcal{D}_{k})}\big{[}\ln\frac{p(x_{t}\>|\>u_{t};\mathcal{D}_{k})}{p(y_{t},x_{t}\>|\>u_{t};\mathcal{D}_{k})}\big{]}=-\frac{1}{2}\ln|R|\,.$ (32)
@@
Let $G_{xx}^{(i)}(\bar{m}_{t})$ be the Hessian of the $i$-th element of the non-linear observation function evaluated at $\bar{m}_{t}$, and let $e_{i}$ be a canonical basis vector. The parameters $\Sigma_{t},\Gamma_{t}$ computed through a second-order Taylor approximation are:
@@
$\mathbb{E}_{p(y_{t},x_{t}\>|\>u_{t};\mathcal{D}_{k})}$ $\big{[}\ln\frac{p(x_{t}\>|\>u_{t};\mathcal{D}_{k})}{p(y_{t},x_{t}\>|\>u_{t};\mathcal{D}_{k})}\big{]}=$ (34)
$-\frac{1}{2}\ln\Big{|}\frac{1}{2}\sum_{i=1}^{D_{y}}\sum_{j=1}^{D_{y}}e_{i}e_{j}^{\intercal}\text{tr}\big{(}G_{xx}^{(i)}(\bar{m}_{t})\bar{S}_{t}G_{xx}^{(j)}(\bar{m}_{t})\bar{S}_{t}\big{)}+R\,\Big{|}\,.$
@@
Interestingly, the ambiguity is constant over $x_{t}$ for the unscented transform.
@@
$\mathbb{E}_{p(y_{t},x_{t}\>|\>u_{t};\mathcal{D}_{k})}$ $\big{[}\ln\frac{p(x_{t}\>|\>u_{t};\mathcal{D}_{k})}{p(y_{t},x_{t}\>|\>u_{t};\mathcal{D}_{k})}\big{]}=-\frac{1}{2}\ln|R|\,.$ (38)
@@ 6 Discussion
One could argue that our analysis is more about model selection than inference, as each Gaussian approximation essentially constitutes a different generative model. In that sense, the experiments only indicate that richer approximations of nonlinear functions lead to better performance, which is not surprising. However, the result is more subtle than that since the unscented transform is richer than the first-order Taylor (produces a more accurate mean estimate [8]) but apparently still leads to constant ambiguity. No, the approximation must be sensitive to how the covariance between states and observations changes as a function of $g$’s curvature. It would be interesting to extend this work with parameter estimation, such as inferring the process noise covariance matrix using a Wishart distribution [15], or the state transition matrix with a Matrix-Normal distribution [1, 12].
@@ Appendix 0.A Appendix: proof of Lemma 1
$=\frac{D_{x}+D_{y}}{2}\ln(2\pi\textit{e})+\frac{1}{2}\ln\big{|}\begin{bmatrix}\bar{S_{t}}&\Gamma_{t}\\ \Gamma^{\intercal}_{t}&\Sigma_{t}\end{bmatrix}\big{|}-\frac{D_{x}}{2}\ln(2\pi\textit{e})+\frac{1}{2}\ln|\bar{S}_{t}|$ (40)
@@ Appendix 0.B Appendix: proof of Theorem 1
$-\frac{1}{2}\ln$ $|\Sigma_{t}-\Gamma_{t}^{\intercal}\bar{S}_{t}^{-1}\Gamma|$ (43)
$=-\frac{1}{2}\ln|G_{x}(\bar{m}_{t})\bar{S}_{t}G_{x}(\bar{m}_{t})^{\intercal}+R-G_{x}(\bar{m}_{t})\bar{S}_{t}^{\intercal}\bar{S}_{t}^{-1}\bar{S}_{t}G_{x}(\bar{m}_{t})^{\intercal}|\,.$
Since the covariance matrix $\bar{S}_{t}$ is symmetric, $\bar{S}_{t}^{\intercal}\bar{S}_{t}^{-1}=\bar{S}_{t}\bar{S}_{t}^{-1}=I$. Thus:
$-\frac{1}{2}\ln|G_{x}(\bar{m}_{t})\bar{S}_{t}G_{x}(\bar{m}_{t})^{\intercal}+R-G_{x}(\bar{m}_{t})\bar{S}_{t}G_{x}(\bar{m}_{t})^{\intercal}|=-\frac{1}{2}\ln|R|\,.$ (44)
@@ Appendix 0.C Appendix: proof of Theorem 2
$-\frac{1}{2}\ln$ $\Big{|}\,\Sigma_{t}-\Gamma_{t}^{\intercal}\bar{S}_{t}^{-1}\Gamma\,\Big{|}=-\frac{1}{2}\ln\Big{|}\,G_{x}(\bar{m}_{t})\bar{S}_{t}G_{x}(\bar{m}_{t})^{\intercal}+$ (45)
$\frac{1}{2}\sum_{i=1}^{D_{y}}\sum_{j=1}^{D_{y}}e_{i}e_{j}^{\intercal}\text{tr}\big{(}G_{xx}^{(i)}(\bar{m}_{t})\bar{S}_{t}G_{xx}^{(j)}(\bar{m}_{t})\bar{S}_{t}\big{)}+R-G_{x}(\bar{m}_{t})\bar{S}_{t}^{\intercal}\bar{S}_{t}^{-1}\bar{S}_{t}G_{x}(\bar{m}_{t})^{\intercal}\Big{|}$
$=$ $-\frac{1}{2}\ln\Big{|}\,\frac{1}{2}\sum_{i=1}^{D_{y}}\sum_{j=1}^{D_{y}}e_{i}e_{j}^{\intercal}\text{tr}\big{(}G_{xx}^{(i)}(\bar{m}_{t})\bar{S}_{t}G_{xx}^{(j)}(\bar{m}_{t})\bar{S}_{t}\big{)}+R\,\Big{|}\,.$ (46)
The covariance matrix $\bar{S}_{t}$ is symmetric. Thus, $\bar{S}_{t}^{\intercal}\bar{S}_{t}^{-1}=\bar{S}_{t}\bar{S}_{t}^{-1}=I$. Note that the Hessian $G_{xx}^{(i)}(\bar{m}_{t})$ depends on the inferred mean of the predicted state $\bar{m}_{t}$.
@@ Appendix 0.D Appendix: proof of Theorem 3
Plugging $\Sigma_{t}$, $\Gamma_{t}$ from (4.3) into the result from Lemma 1, gives:
$-\frac{1}{2}\ln\big{|}\Sigma_{t}-\Gamma_{t}^{\intercal}\bar{S}_{t}^{-1}\Gamma\big{|}=-\frac{1}{2}\ln\Big{|}\sum_{i^{\prime}=0}^{2D_{x}}w_{i^{\prime}}(g(\chi_{i^{\prime}})\!-\!\mu_{t})(g(\chi_{i^{\prime}})\!-\!\mu_{t})^{\intercal}+R$
@@
Column selection $\big{[}\cdot\big{]}_{i}$ is equivalent to right-multiplication with a canonical basis vector $e_{i}$;
@@
$\sqrt{\bar{S}_{t}}^{\intercal}\bar{S}_{t}^{-1}\sqrt{\bar{S}_{t}}$ $=\big{(}V\Omega^{1/2}V^{-1}\big{)}^{\intercal}(V\Omega V^{-1})^{-1}\big{(}V\Omega^{1/2}V^{-1}\big{)}$ (53)
@@
$-\frac{1}{2}\ln|\sum_{i^{\prime}=0}^{2D_{x}}w_{i^{\prime}}(g(\chi_{i^{\prime}})\!-\!\mu_{t})(g(\chi_{i^{\prime}})\!-\!\mu_{t})^{\intercal}\!+\!R\!-\!\sum_{i=0}^{2D_{x}}w_{i}$ $(g(\chi_{i})\!-\!\mu_{t})(g(\chi_{i})\!-\!\mu_{t})^{\intercal}|$
$=-\frac{1}{2}\ln|R|\,.$ (59)
This concludes the proof.

### v2

# TU Eindhoven, Eindhoven, Netherlands

@@ 1 Introduction
Active inference agents are based on free energy functionals that rank policies on explorative and goal-directed behaviour [5, 7, 6, 16]. The expected free energy functional can be understood through its decomposition into a cross-entropy term between states and observations given action ("ambiguity"), and a Kullback-Leibler divergence between the posterior predictive and a goal prior distribution ("risk") [7, 18, 4]. We show that Gaussian approximations of a non-linear observation function that are itself linear in the covariance matrix, e.g., first-order Taylor and the unscented transform [9], lead to ambiguity terms that are constant over states. This echoes an earlier finding that agents with a linear Gaussian state-space model exhibit a constant ambiguity term [10]. However, utilizing a second-order Taylor approximation induces a non-constant ambiguity term. Under this model, the agent will avoid states where the non-linear measurement function curves strongly. Our contributions are:
@@ 2 Problem statement
We want to plan a trajectory for a robot across a plane. The robot’s state at time $k$ is its planar position and time derivatives, $x_{k}\in\mathbb{R}^{D_{x}}$. The robot does not sense position directly, but has to infer it from noisy measurements $y_{k}\in\mathbb{R}^{D_{y}}$, produced by a sensor through a non-linear mapping $g:\mathbb{R}^{D_{x}}\rightarrow\mathbb{R}^{D_{y}}$ and measurement noise $v_{k}\in\mathcal{R}^{D_{y}}$. It accepts control inputs $u_{k}\in\mathbb{R}^{D_{u}}$ and moves according to linear dynamics with a transition matrix $A\in\mathbb{R}^{D_{x}\times D_{x}}$, control matrix $B\in\mathbb{R}^{D_{x}\times D_{u}}$ and process noise $e_{k}\in\mathbb{R}^{D_{x}}$. Overall, we consider robot systems described with discrete-time state-space models of the form:
@@
where $Q,R$ are noise covariance matrices.
@@ 2.0.1 Example
$g(x_{k})=\begin{bmatrix}\phi_{k}\\ d_{k}\end{bmatrix}=\begin{bmatrix}\sqrt{x_{1k}^{2}+x_{2k}^{2}}\\ \arctan(x_{1k},x_{2k})\end{bmatrix}\,,\ R=\begin{bmatrix}\rho^{2}_{1}&0\\ 0&\rho^{2}_{2}\end{bmatrix}\,,$ (4)
@@ 3.1 Probabilistic Model
This distribution is linear in $x_{k}$, and will allow for exact Bayesian filtering. But note that the parameters $\mu_{k}$, $\Gamma_{k}$ and $\Sigma_{k}$ may be nonlinear functions of $x_{k}$, depending on the type of Gaussian approximation (specifics treated in Section 4), and may thus capture more of the effect of $g(x_{k})$.
@@ 3.3 Inferring controls
$q(y_{t}\>|\>x_{t})\triangleq p(y_{t}\>|\>x_{t})\,,\quad q(x_{t},u_{t})\triangleq p(x_{t}\>|\>u_{t};\mathcal{D}_{k})q(u_{t})\,.$ (14)
@@
"Risk" refers to the Kullback-Leibler (KL) divergence between predicted and desired future observations. The inner integral in the risk term leads to a Gaussian distribution (Eq. 24) and the KL divergence between Gaussians is [3]:
@@
where $\Psi_{*}\triangleq(\mu_{*}-\mu_{t}\big{)}\big{(}\mu_{*}-\mu_{t})^{\intercal}$. "Ambiguity" refers to the conditional entropy of the future observations given the future states.
@@
$\mathcal{J}_{k}(u_{t})\!=\!\frac{1}{2}\Big{(}\ln\frac{|\Sigma_{*}|}{|\Sigma_{t}|}\!+\!D_{y}\ln(2\pi)\!+\!\text{tr}\big{(}\Sigma_{*}^{-1}(\Sigma_{t}\!+\!\Psi_{*})\big{)}\!+\!\ln\big{|}\Sigma_{t}\!-\!\Gamma^{\intercal}_{t}\bar{S}_{t}^{\text{-}1}\Gamma_{t}\big{|}\Big{)}\,.$ (29)
Note that $\Gamma_{t}$, $\Sigma_{t}$ and $\Psi_{*}$ depend on $u_{t}$. The above steps can be generalized to a longer time horizon $t=k+1,\dots k+T$. Because the prior is independent over time, $p(\bar{u})=\prod_{t=1}^{T}p(u_{t})$ (see Eq. 6), the function $\mathcal{J}_{k}(\bar{u})$ factorizes to a sum of recursive expected free energy functions $\sum_{t=1}^{T}\mathcal{J}_{k}(u_{t})$.
@@ 4 Gaussian approximations
$\mathbb{E}_{p(y_{t},x_{t}\>|\>u_{t};\mathcal{D}_{k})}\big{[}-\ln\frac{p(y_{t},x_{t}\>|\>u_{t};\mathcal{D}_{k})}{p(x_{t}\>|\>u_{t};\mathcal{D}_{k})}\big{]}=\frac{D_{y}}{2}\ln(2\pi\textit{e})+\frac{1}{2}\ln|R|\,.$ (32)
@@
Let $G_{xx}^{(i)}(\bar{m}_{t})$ be the Hessian of the $i$-th element of the non-linear observation function evaluated at $\bar{m}_{t}$ and $e_{i}$ be a canonical basis vector. The parameters $\Sigma_{t},\Gamma_{t}$ computed through a second-order Taylor approximation are:
@@
$\mathbb{E}_{p(y_{t},x_{t}\>|\>u_{t};\mathcal{D}_{k})}\big{[}-\ln\frac{p(y_{t},x_{t}\>|\>u_{t};\mathcal{D}_{k})}{p(x_{t}\>|\>u_{t};\mathcal{D}_{k})}\big{]}=$ (34)
$\qquad\frac{D_{y}}{2}\ln(2\pi\textit{e})+\frac{1}{2}\ln\Big{|}\frac{1}{2}\sum_{i=1}^{D_{y}}\sum_{j=1}^{D_{y}}e_{i}e_{j}^{\intercal}\text{tr}\big{(}G_{xx}^{(i)}(\bar{m}_{t})\bar{S}_{t}G_{xx}^{(j)}(\bar{m}_{t})\bar{S}_{t}\big{)}+R\,\Big{|}\,.$
@@
Interestingly, the ambiguity is also constant for the unscented transform.
@@
$\mathbb{E}_{p(y_{t},x_{t}\>|\>u_{t};\mathcal{D}_{k})}$ $\big{[}-\ln\frac{p(y_{t},x_{t}\>|\>u_{t};\mathcal{D}_{k})}{p(x_{t}\>|\>u_{t};\mathcal{D}_{k})}\big{]}=\frac{D_{y}}{2}\ln(2\pi\textit{e})+\frac{1}{2}\ln|R|\,.$ (38)
@@ 6 Discussion
One could argue that our analysis is more about model selection than inference, as each Gaussian approximation essentially constitutes a different generative model. In that sense, the experiments only indicate that richer approximations of nonlinear functions lead to better performance, which is not surprising. However, the result is more subtle than that since the unscented transform is richer than the first-order Taylor (produces a more accurate mean estimate [8]) but apparently still leads to constant ambiguity. No, the approximation must be sensitive to how the covariance matrix of the joint distribution over states and observations changes as a function of $g$’s curvature. It would be interesting to extend this work with parameter estimation, such as inferring the process noise covariance matrix using a Wishart distribution [15], or the state transition matrix with a Matrix-Normal distribution [1, 12].
@@ Appendix 0.A Appendix: proof of Lemma 1
$=\frac{D_{x}+D_{y}}{2}\ln(2\pi\textit{e})+\frac{1}{2}\ln\big{|}\begin{bmatrix}\bar{S_{t}}&\Gamma_{t}\\ \Gamma^{\intercal}_{t}&\Sigma_{t}\end{bmatrix}\big{|}-\Big{(}\frac{D_{x}}{2}\ln(2\pi\textit{e})+\frac{1}{2}\ln|\bar{S}_{t}|\Big{)}$ (40)
@@ Appendix 0.B Appendix: proof of Theorem 1
$\frac{D_{y}}{2}\ln(2\pi\textit{e})+\frac{1}{2}\ln|\Sigma_{t}-\Gamma_{t}^{\intercal}\bar{S}_{t}^{-1}\Gamma|$ (43)
$\quad=\frac{D_{y}}{2}\ln(2\pi\textit{e})+\frac{1}{2}\ln|G_{x}(\bar{m}_{t})\bar{S}_{t}G_{x}(\bar{m}_{t})^{\intercal}+R-G_{x}(\bar{m}_{t})\bar{S}_{t}^{\intercal}\bar{S}_{t}^{-1}\bar{S}_{t}G_{x}(\bar{m}_{t})^{\intercal}|$
$\quad=\frac{D_{y}}{2}\ln(2\pi\textit{e})+\frac{1}{2}\ln|R|\,.$ (44)
The cancellation is due to $\bar{S}_{t}$ being symmetric, i.e., $\bar{S}_{t}^{\intercal}\bar{S}_{t}^{-1}=\bar{S}_{t}\bar{S}_{t}^{-1}=I$.
@@ Appendix 0.C Appendix: proof of Theorem 2
$\frac{D_{y}}{2}\ln(2\pi\textit{e})\!+\!\frac{1}{2}\ln\Big{|}\,\Sigma_{t}\!-\!\Gamma_{t}^{\intercal}\bar{S}_{t}^{-1}\Gamma\,\Big{|}=\!\frac{D_{y}}{2}\ln(2\pi\textit{e})\!+\!\frac{1}{2}\ln\Big{|}\,G_{x}(\bar{m}_{t})\bar{S}_{t}G_{x}(\bar{m}_{t})^{\intercal}$ (45)
$\quad+\frac{1}{2}\sum_{i=1}^{D_{y}}\sum_{j=1}^{D_{y}}e_{i}e_{j}^{\intercal}\text{tr}\big{(}G_{xx}^{(i)}(\bar{m}_{t})\bar{S}_{t}G_{xx}^{(j)}(\bar{m}_{t})\bar{S}_{t}\big{)}+R-G_{x}(\bar{m}_{t})\bar{S}_{t}^{\intercal}\bar{S}_{t}^{-1}\bar{S}_{t}G_{x}(\bar{m}_{t})^{\intercal}\Big{|}$
$\ =\frac{D_{y}}{2}\ln(2\pi\textit{e})+\frac{1}{2}\ln\Big{|}\,\frac{1}{2}\sum_{i=1}^{D_{y}}\sum_{j=1}^{D_{y}}e_{i}e_{j}^{\intercal}\text{tr}\big{(}G_{xx}^{(i)}(\bar{m}_{t})\bar{S}_{t}G_{xx}^{(j)}(\bar{m}_{t})\bar{S}_{t}\big{)}+R\,\Big{|}\,.$ (46)
The covariance matrix $\bar{S}_{t}$ is symmetric, i.e., $\bar{S}_{t}^{\intercal}\bar{S}_{t}^{-1}=\bar{S}_{t}\bar{S}_{t}^{-1}=I$. Note that the Hessian $G_{xx}^{(i)}(\bar{m}_{t})$ depends on the inferred mean of the predicted state $\bar{m}_{t}$, meaning that ambiguity is not constant over state-space.
@@ Appendix 0.D Appendix: proof of Theorem 3
Plugging $\mu_{t}$, $\Sigma_{t}$, $\Gamma_{t}$ from (4.3) into the log-determinant term from the result in Lemma 1, gives:
$\frac{1}{2}\ln\big{|}\Sigma_{t}-\Gamma_{t}^{\intercal}\bar{S}_{t}^{-1}\Gamma\big{|}=\frac{1}{2}\ln\Big{|}\sum_{i^{\prime}=0}^{2D_{x}}w_{i^{\prime}}(g(\chi_{i^{\prime}})\!-\!\mu_{t})(g(\chi_{i^{\prime}})\!-\!\mu_{t})^{\intercal}+R$
@@
Furthermore, note that column selection $\big{[}\cdot\big{]}_{i}$ is equivalent to right-multiplication with a canonical basis vector $e_{i}$;
@@
$\sqrt{\bar{S}_{t}}^{\intercal}\bar{S}_{t}^{-1}\sqrt{\bar{S}_{t}}$ $=\big{(}V\Omega^{1/2}V^{-1}\big{)}^{\intercal}V\Omega^{-1}V^{-1}\big{(}V\Omega^{1/2}V^{-1}\big{)}$ (53)
@@
$\frac{1}{2}\ln|\sum_{i^{\prime}=0}^{2D_{x}}w_{i^{\prime}}(g(\chi_{i^{\prime}})\!-\!\mu_{t})(g(\chi_{i^{\prime}})\!-\!\mu_{t})^{\intercal}\!+\!R\!-\!\sum_{i=0}^{2D_{x}}w_{i}$ $(g(\chi_{i})\!-\!\mu_{t})(g(\chi_{i})\!-\!\mu_{t})^{\intercal}|$
$=\frac{1}{2}\ln|R|\,.$ (59)
Using this result and Lemma 1, we have proven Theorem 3.